1. A discount store sold plastic cups for $3.25 each and ceramic cups for $4.50 each.
a. If 500 cups were sold for a total of $2,090, how many cups of each type were sold?

Answer by checkley77(12569)

4.50X+3.25(500-X)=2,090
4.50X+1,625-3.25X=2,090
1.25X=2,090-1,625
1.25X=465
X=465/1.25
X=372 ANS. FOR THE NUMBER OF CERAMIC CUPS.
500-372=128 ANS. COST OF THE NUMBER OF PLASTIC CUPS.
PROOF:
4.50*372+3.25*128=2,090
1,674+416=2,090
2,090=2,090